Every cereal box has a gift inside, but you cannot tell from the outside what the gift is. The store manager assures you that 18 of the 51 boxes on the shelf have the secret decoder ring. The other 33 boxes on the shelf have a different gift inside. If you randomly select two boxes of cereal from the shelf to purchase, what is the probability that BOTH of them have the secret decoder ring?

(Give answer as a decimal correct to four decimal places.)
Answer #1

Answer:

Given,

To give the probability that BOTH of them have the secret decoder ring

Required probability = P(1st has the secret decoder ring)*P(2nd has given 1st)

= (18/51) * (17/50)

= 0.35294*0.34

= 0.1200

So the probability that BOTH of them have the secret decoder ring is 0.1200